# Conformal Validity Guarantees Exist for Any Data Distribution

### v2

@@ C.2 The Role of Weighted Exchangeability in Conformal Prediction
This section sketches the argument in appendix A.7 of Tibshirani et al. (2019), which proves a key lemma underlying the validity of weighted CP, with the goal of understanding the role of the weight functions in an inversion of a weighted permutation test.
@@
We can then write out this statement more explicitly, as we did before, by including the random variables in the weight and factor function arguments:
$\mathbb{P}\{V_{n+1}=v_{i}\ |\ E_{z}\}=$ $\frac{\sum_{\sigma:\sigma(n+1)=i}\prod_{j=1}^{n+1}w_{j}(Z_{j}=z_{\sigma(j)})g(Z_{1}=z_{\sigma(1)},...,Z_{n+1}=z_{\sigma(n+1)})}{\sum_{\sigma}\prod_{j=1}^{n+1}w_{j}(Z_{j}=z_{\sigma(j)})g(Z_{1}=z_{\sigma(1)},...,Z_{n+1}=z_{\sigma(n+1)})}.$
In this more explicit version of the statement, we see that $w_{j}$ (for each $j\in\{1,...,n+1\}$) does not only take as its argument the value $z_{\sigma(j)}$, but rather it is a function of the event $Z_{j}=z_{\sigma(j)}$. (The reliance of $w_{j}$ on $Z_{j}$ in the original notation is implicit in the index $j$.) Our more explicit notation can thus allow us to see that—so long as the data are weighted exchangeable—the scalar quantity $\prod_{j=1}^{n+1}w_{j}(Z_{j}=z_{\sigma(j)})$ can be thought of as how much more likely it would be to observe $Z_{1}=z_{\sigma(1)},...,Z_{n+1}=z_{\sigma(n+1)}$ (given $E_{z}$), relative to the measure that our core function $g$ assigns to this event—i.e., relative to $g(Z_{1}=z_{\sigma(1)},...,Z_{n+1}=z_{\sigma(n+1)})$. Put more simply, $\prod_{j=1}^{n+1}w_{j}(Z_{j}=z_{\sigma(j)})$ tells us how much we should “weigh” or “count” a given permutation $\sigma$.
Simplifying using the fact that $g(Z_{1}=z_{\sigma(1)},...,Z_{n+1}=z_{\sigma(n+1)})=g(Z_{1}=z_{1},...,Z_{n+1}=z_{n+1})$ under weighted exchangeability yields
$\mathbb{P}\{V_{n+1}=v_{i}\ |\ E_{z}\}=$ $\frac{\sum_{\sigma:\sigma(n+1)=i}\prod_{j=1}^{n+1}w_{j}(Z_{j}=z_{\sigma(j)})g(Z_{1}=z_{1},...,Z_{n+1}=z_{n+1})}{\sum_{\sigma}\prod_{j=1}^{n+1}w_{j}(Z_{j}=z_{\sigma(j)})g(Z_{1}=z_{1},...,Z_{n+1}=z_{n+1})}$
$=$ $\frac{\sum_{\sigma:\sigma(n+1)=i}\prod_{j=1}^{n+1}w_{j}(Z_{j}=z_{\sigma(j)})}{\sum_{\sigma}\prod_{j=1}^{n+1}w_{j}(Z_{j}=z_{\sigma(j)})}.$
This last line shows explicitly that $\mathbb{P}\{V_{n+1}=v_{i}\ |\ E_{z}\}$ is given by the total “weight” of the permutations $\sigma$ such that $\sigma(n+1)=i$, normalized over the total weight of all possible permutations. (We stop here because our emphasis is on the role of the weight functions; the rest of the argument follows similarly as in the previous subsection.)
To review: For each $j\in\{1,...,n+1\}$, the weight function quantity $w_{j}(z_{\sigma(j)})=w_{j}(Z_{j}=z_{\sigma(j)})$ weighs the relative likelihood or “importance” of the event $Z_{j}=z_{\sigma(j)}$, while the product $\prod_{j=1}^{n+1}w_{j}(Z_{j}=z_{\sigma(j)})$ weighs the the relative likelihood or “importance” of the joint event of observing all the data in a particular ordering or permutation $\sigma$ (given $E_{z}$, the observed set of data values), that is of observing $Z_{1}=z_{\sigma(1)},...,Z_{n+1}=z_{\sigma(n+1)}$. So, the product $\prod_{j=1}^{n+1}w_{j}(Z_{j}=z_{\sigma(j)})$ thus tells us how much to “count” a given permutation $\sigma$.
C.3 The Role of Pseudo-Exchangeability in Conformal Prediction
@@
Here, in the same way that for weighted exchangeable data the product $\prod_{j=1}^{n+1}w_{j}(z_{\sigma(j)})$ represents the (relative) likelihood of observing the joint sequence of events $Z_{1}=z_{\sigma(1)},...,Z_{n+1}=z_{\sigma(n+1)}$, for pseudo-exchangeable data the analogous role is played by the product $\prod_{j=1}^{n+1}g_{j}(z_{\sigma(j)};z_{-\sigma(j)})$: the role of this product is to tell us how much to “weigh” or “count” a given permutation $\sigma$. However, the specification in pseudo-exchangeability’s definition that “$g_{i}(\cdot;v_{-i})$ is a function that depends on the multiset $v_{-i}$ (that is, on the values in $v_{-i}$ but not on their ordering)” (emphasis added) prevents pseudo-exchangeability from formally describing multistep feedback covariate shift. We state this more formally in the following remark.
Remark C.1.
@@
Pseudo-exchangeability does not formally characterize multistep FCS. Consider a two-step instance of FCS for a counterexample, that is with data $Z_{1},...,Z_{n+2}$ generated under MFCS. The factor function for the first step $g_{n+1}(X_{n+1}=x_{n+1}\ ;\ Z_{1}=z_{1},...,Z_{n}=z_{n})$ depends on $Z_{1},...,Z_{n}$ differently than it does on $Z_{n+2}$ (i.e., it does not depend on $Z_{n+2}$ at all). So, the factor function $g_{n+1}(\cdot\ ;\ Z_{1}=z_{1},...,Z_{n}=z_{n})$ cannot be said to depend on the multiset of values $z_{-(n+1)}=\{z_{1},...,z_{n},z_{n+2}\}$ independently of the ordering of the values (i.e., it requires knowing which of the values corresponds to $Z_{1:n}$ versus $Z_{n+2}$), as is required by the defintion of pseudo-exchangeability from Fannjiang et al. (2022).
Appendix D Additonal Experimental Details
D.1 Black-box Optimization Experimental Details
@@
Protein Design Datasets Our multistep biomolecular design experiments leverage fluorescent protein data from Poelwijk et al. (2019), previously used for evaluation of related single-step FCS CP methods in Fannjiang et al. (2022) and Prinster et al. (2023). This dataset has the benefit of a combinatorially complete set of labels, which simulates measuring a design in an experiment. In particular, both a “blue” and “red” wavelength fluorescence strength were experimentally measured for each of $1922^{13}=8,$ 192 possible combinations of binary variations to a wild-type (natural) fluorescent protein at 13 positions on the primary sequence. As in Fannjiang et al. (2022), we set the noise scale for each sequence $i$ to the residual norm between that sequence’s label and a separately fit linear model.
D.2 Active Learning Experimental Details
@@
Datasets for Active Learning Experiments We evaluate on four datasets from the UCI Machine Learning Repository (Frank, 2010) that are commonly used for evaluation in the CP literature (Tibshirani et al., 2019; Barber et al., 2021; Prinster et al., 2022, 2023): the NASA airfoil self-noise dataset (1503 samples, $p=5$) (Brooks et al., 2014), the communities and crime dataset (1994 samples, $p=99$) (Redmond, 2009), the Medical Expenditure Panel Survey 2016 data set [^18] https://meps.ahrq.gov/mepsweb/data_stats/download_data_files_detail.jsp?cboPufNumber=HC-192 (33005 samples, $p=107$) (preprocessed as in Barber et al. (2021); details for an older version of the data are in Ezzati-Rice et al. (2008)), and the blog feedback dataset (60021 samples, $p=281$) (Buza, 2014).
@@
Model for Active Learning Experiments: The model we used for the active learning experiments is scikit-learn’s GaussianProcessRegressor, with kernel=DotProduct()+WhiteKernel() and default parameters otherwise.
Appendix E Additional Experimental Results
E.1 Bounding the Query Function to Enforce Informative (Finite) Prediction Intervals
Because noninformative prediction sets ($\widehat{\mathcal{C}}_{n}(x)=\mathcal{Y}$) occur in weighted CP when a test point’s normalized weight as in Eq. (9) exceeds $\alpha$, we now consider redesigning the query probability functions to prevent this from happening. For simplicity, we impose this constraint on the one-step weights. In particular, we consider bounding the query probability function $\mathbb{P}(x\mid Z_{1},...,Z_{j-1})$ so that instead of being proportional to $\exp(\lambda\cdot u_{t}(x))$, the probabilities are instead proportional to a bounded function:
$\mathbb{P}(x\mid Z_{1},...,Z_{j-1})\propto$ $\min\big{(}\exp(\lambda\cdot u_{t}(x)),B\big{)},$
where $\text{where }B=\sup\bigg{\{}b>0:\frac{b}{\sum_{i=1}^{n}\min\big{(}\exp(\lambda\cdot u_{t}(x)(X_{i})),b\big{)}+b}<\alpha\bigg{\}}.$
That is, we select $B$ as the largest bound that we could use to avoid infinite prediction intervals (in the depth $d=1$ weights). This is because
$\min\big{(}\exp(\lambda\cdot u_{t}(x)),B\big{)}\leq B$
Across the four real-world datasets and the same experimental conditions we used for our active learning experiments in the main paper Figure 3, the following results empirically validate that imposing the proposed adaptive bound on our MFCS CP method’s query function completely prevents any noninformative (infinite width) intervals, all without sacrificing coverage even over a long time horizon (70 active learning steps). For comparison, for each dataset we provide the corresponding results for the original unbounded query function. We moreover plot the “relative magnitude” of the bound $B$—i.e., we plot $B/B_{\max}$, where $B_{\max}=\max_{x\in\text{pool}}(\exp(\lambda\cdot\hat{\sigma}(x)))$—across the same trajectory of active learning steps.
Airfoil dataset
Figure 4: Unbounded (top row) versus bounded (bottom row) active learning experiments of proposed multistep split CP method for $d=3$ (red circles) compared to baselines of unweighted split CP (orange squares), one-step split CP (green triangles), and ACI (gray squares) on the airfoil dataset. The Y-axes represent mean coverage, median interval width, and mean squared error on a holdout test set; the X-axes correspond to the number of active learning query steps, with each query based on posterior variance of a GP regressor. All values are computed over 350 distinct random seeds.
Communities dataset
Figure 5: Unbounded (top row) versus bounded (bottom row) active learning experiments of proposed multistep split CP method for $d=3$ (red circles) compared to baselines of unweighted split CP (orange squares), one-step split CP (green triangles), and ACI (gray squares) on the communities dataset. The Y-axes represent mean coverage, median interval width, and mean squared error on a holdout test set; the X-axes correspond to the number of active learning query steps, with each query based on posterior variance of a GP regressor. All values are computed over 350 distinct random seeds.
Medical Expenditure Panel Survey (MEPS) dataset
Figure 6: Unbounded (top row) versus bounded (bottom row) active learning experiments with proposed multistep split CP method for estimation depth $d=3$ (red circles) compared to baselines of unweighted split CP (orange squares), one-step split CP (green triangles), and ACI (gray squares) on the Medical Expenditure Panel Survey (MEPS) dataset. The Y-axes represent mean coverage, median interval width, mean squared error on a holdout test set, and the relative bound magnitude $B/B_{\text{max}}$; the X-axes correspond to the number of active learning query steps, with each query based on posterior variance of a GP regressor. All values are computed over 350 distinct random seeds.
Blog dataset
Figure 7: Unbounded (top row) versus bounded (bottom row) active learning experiments of proposed multistep split CP method for $d=3$ (red circles) compared to baselines of unweighted split CP (orange squares), one-step split CP (green triangles), and ACI (gray squares) on the blog dataset. The Y-axes represent mean coverage, median interval width, and mean squared error on a holdout test set; the X-axes correspond to the number of active learning query steps, with each query based on posterior variance of a GP regressor. All values are computed over 350 distinct random seeds.
The bounded query function’s improvement in prediction-interval informativeness sometimes comes at the cost of an initially delayed decrease in MSE—where the constraint forces the agent to initially be more “cautious” in its exploration—though the long-run MSE for the bounded and unbounded query functions are comparable for most datasets. The increasing bound magnitude over time demonstrates that this bounded query function enables the active-learning agent to be less cautious over time by relaxing the constraint of the bound once it has seen more data.

### v3

@@ D.2 The Role of Weighted Exchangeability in Conformal Prediction
We now sketch the argument in appendix A.7 of Tibshirani et al. (2019), which proves a key lemma underlying the validity of weighted CP, with the goal of understanding the role of the weight functions in an inversion of a weighted permutation test.
@@
Recalling the definition of weighted exchangeability wherein $g$ is a permutation-invariant function, we are able to substitute $g(z_{\sigma(1)},...,z_{\sigma(n+1)})=g(z_{1},...,z_{n+1})$, analogously to how we substituted $f(z_{\sigma(1)},...,z_{\sigma(n+1)})=f(z_{1},...,z_{n+1})$ using exchangeability in Appendix D.1:
$\mathbb{P}\{V_{n+1}=v_{i}\ |\ E_{z}\}=$ $\frac{\sum_{\sigma:\sigma(n+1)=i}\prod_{j=1}^{n+1}w_{j}(z_{\sigma(j)})g(z_{1},...,z_{n+1})}{\sum_{\sigma}\prod_{j=1}^{n+1}w_{j}(z_{\sigma(j)})g(z_{1},...,z_{n+1})}$
$=$ $\frac{\sum_{\sigma:\sigma(n+1)=i}\prod_{j=1}^{n+1}w_{j}(z_{\sigma(j)})}{\sum_{\sigma}\prod_{j=1}^{n+1}w_{j}(z_{\sigma(j)})}.$
From this simplified statement, it is relatively easier to see that the product $\prod_{j=1}^{n+1}w_{j}(z_{\sigma(j)})$ can be thought of as how much more likely it would be to observe $Z_{1}=z_{\sigma(1)},...,Z_{n+1}=z_{\sigma(n+1)}$ (given $E_{z}$), relative to the measure that our core function $g$ assigns to this event—i.e., relative to $g(z_{\sigma(1)},...,z_{\sigma(n+1)})=g(z_{1},...,z_{n+1})$. Put more simply, $\prod_{j=1}^{n+1}w_{j}(z_{\sigma(j)})$ tells us how much we should “weigh” or “count” a given permutation $\sigma$.
D.3 The Role of Pseudo-Exchangeability in Conformal Prediction
@@
Here, in the same way that for weighted exchangeable data the product $\prod_{j=1}^{n+1}w_{j}(z_{\sigma(j)})$ represents the (relative) likelihood of observing the joint sequence of events $Z_{1}=z_{\sigma(1)},...,Z_{n+1}=z_{\sigma(n+1)}$ (given $E_{Z}$), for pseudo-exchangeable data the analogous role is played by the product $\prod_{j=1}^{n+1}g_{j}(z_{\sigma(j)};z_{-\sigma(j)})$: the role of this product is to tell us how much to “weigh” or “count” a given permutation $\sigma$. However, the specification in pseudo-exchangeability’s definition that “$g_{i}(\cdot;v_{-i})$ is a function that depends on the multiset $v_{-i}$ (that is, on the values in $v_{-i}$ but not on their ordering)” (emphasis added) prevents pseudo-exchangeability from formally describing multistep feedback covariate shift. We state this more formally in the following remark.
Remark D.1.
@@
Pseudo-exchangeability does not formally characterize multistep FCS. Consider a two-step instance of FCS for a counterexample, that is with data $Z_{1},...,Z_{n+2}$ generated under MFCS. The factor function for the first step $g_{n+1}(z_{n+1}\ ;\ Z_{1}=z_{1},...,Z_{n}=z_{n})$ depends on $Z_{1},...,Z_{n}$ differently than it does on $Z_{n+2}$ (i.e., it does not depend on $Z_{n+2}$ at all). So, the factor function $g_{n+1}(\cdot\ ;\ Z_{1}=z_{1},...,Z_{n}=z_{n})$ cannot be said to depend on the multiset of values $z_{-(n+1)}=\{z_{1},...,z_{n},z_{n+2}\}$ independently of the ordering of the values (i.e., it requires knowing which of the values corresponds to $Z_{1:n}$ versus $Z_{n+2}$), as is required by the defintion of pseudo-exchangeability from Fannjiang et al. (2022).
Appendix E Additonal Details for Main Paper Experiments
E.1 Black-box Optimization Experimental Details
@@
Protein Design Datasets Our multistep biomolecular design experiments leverage fluorescent protein data from Poelwijk et al. (2019), previously used for evaluation of related single-step FCS CP methods in Fannjiang et al. (2022) and Prinster et al. (2023). This dataset has the benefit of a combinatorially complete set of labels, which simulates measuring a design in an experiment. In particular, both a “blue” and “red” wavelength fluorescence strength were experimentally measured for each of $1922^{13}=8,$ 192 possible combinations of binary variations to a wild-type (natural) fluorescent protein at 13 positions on the primary sequence.
Measurement Noise for Protein Design Experiments We followed the same procedure for adding measurement noise as described in Fannjiang et al. (2022)’s Supplementary Materials S3. That is, every time the $i$-th sequence was sampled from the combinatorially complete library (including for initial training and calibration data as well as for each query point), zero-mean Guassian noise was added (to the label value) with standard deviation set to the absolute residual between that $i$-th sequence’s true label and the prediction on that $i$-th sequence from a separately fit linear model. The separately-fit linear model had as its input covariates the up to seventh-order terms identified as statistically significant by Poelwijk et al. (2019), as described by Fannjiang et al. (2022).
Table 1: Hyperparameters for Full CP protein design experiments with ridge regression predictor (Figure 1).
Hyperparameter name Value(s)
$\alpha$ (Target miscoverage rate) 0.1
$n$ (Initial number of training/calibration samples) 32
$\lambda$ (Shift magnitude or “inverse temperature”) 8.0
$d$ (MFCS estimation depth) $\{1,2\}$
$T$ (Number of MFCS steps past IID initialization) 5
Ridge regression regularization strength 0.01
Random seeds $\{0,...,999\}$
Table 2: Hyperparameters for Split CP protein design neural network experiments (Figure 2) with the MLPRegressor from scikit-learn (L-BFGS solver and logistic activation function, default parameters otherwise).
Hyperparameter name Value(s)
$\alpha$ (Target miscoverage rate) 0.1
$n_{\text{train}}$ (Initial number of training samples) 32
$n_{\text{cal}}$ (Initial number of calibration samples) 32
$\lambda$ (Shift magnitude or “inverse temperature”) 5.0
$d$ (MFCS estimation depth) $\{1,2,3,4\}$
$T$ (Number of MFCS steps past IID initialization) 10
Probability each queried point is added to training (versus calibration) data 0.5
Random seeds $\{0,...,499\}$
E.2 Active Learning Experimental Details
@@
Datasets for Active Learning Experiments We evaluated on four datasets from the UCI Machine Learning Repository (Frank, 2010) that are commonly used for evaluation in the CP literature (Tibshirani et al., 2019; Barber et al., 2021; Prinster et al., 2022, 2023) and represent a range of sample sizes and dimensionalities: the NASA airfoil self-noise dataset (1503 samples, $p=5$) (Brooks et al., 2014), the communities and crime dataset (1994 samples, $p=99$) (Redmond, 2009), the Medical Expenditure Panel Survey 2016 data set [^23] https://meps.ahrq.gov/mepsweb/data_stats/download_data_files_detail.jsp?cboPufNumber=HC-192 (33005 samples, $p=107$) (preprocessed as in Barber et al. (2021); details for an older version of the data are in Ezzati-Rice et al. (2008)), and the blog feedback dataset (60021 samples, $p=281$) (Buza, 2014).
Sampling Holdout Test Set and Initial Training Set For all experiments, a holdout test set of 250 samples was first sampled uniformly at random from the full dataset to track the accuracy of the base predictor as the active learning progressed. To simulate a practical active learning setting where there often tends to be sample-selection bias in how the initial training data is obtained relative to the desired test distribution, we sampled the initial training and calibration datasets from the remaining non-test-set data with probabilities proportional to $\exp(\gamma\cdot X_{\text{PCA}1}^{\text{normed}})$, where $X_{\text{PCA}1}^{\text{normed}}$ is the min-max-normed first principal component representation of the data (that is, the dataset projected onto the first principal component extracted from the same dataset).
Measurement Noise for Active Learning Experiments Every time the $i$-th datapoint was sampled (including for the holdout test set, initial training/calibration data, and each queried point), zero-mean Gaussian noise was added to the label value. The standard deviation for the Gaussian noise for the $i$-th datapoint was set as 0.05 times the absolute residual between that $i$-th point’s true label value and the prediction for that $i$-th point by a separately fit kernel ridge regression model that was fit to all of the data samples. In particular, the separately-fit regressor was the KernelRidge method from scikit-learn, with regularization strength 1.0 and default parameters otherwise.
@@
Table 3: Hyperparameters for Split CP active learning experiments (Figure 3) with the GaussianProcessRegressor from scikit-learn as the active learning agent (kernel=DotProduct(sigma_0=0.05)+WhiteKernel(noise_level=0.05), default parameters otherwise).
Hyperparameter name Value(s)
$\alpha$ (Target miscoverage rate) 0.1
$n_{\text{train}}$ (Initial number of training samples) 64
$n_{\text{cal}}$ (Initial number of calibration samples) 16
$\lambda$ (Shift magnitude) 10.0
$d$ (MFCS estimation depth) $\{1,3\}$
$T$ (Number of MFCS steps past IID initialization) 70
$\gamma$ (Bias magnitude in initial training data IID sampling) 3.0
ACI baseline “step size” (default parameter in Gibbs & Candès (2021)) 0.005
Probability each queried point is added to training (versus calibration) data 0.5
Random seeds $\{0,...,349\}$
Appendix F Additional Experiments
F.1 Shift Magnitude Ablation Study
These ablation studies examine the effect of changing the shift magnitude $\lambda$, or the “aggressiveness” of the ML agent’s query strategy, focusing on the Split CP blue fluorescent protein design case. Other than the different values of $\lambda$, the experimental setting and other hyperparameters are the same as in the blue protein results reported in the main paper Figure 2, where $\lambda=5.0$ (full hyperparameters reported in Appendix E.1, Table 2). Increasing $\lambda$ tends to exacerbate the discrepancy in coverage between the $\{1,2,3,4\}$-step MFCS CP methods (with higher-order methods maintaining stronger coverage); increase interval widths of the MFCS CP methods; and increase the fitness values of the designed protein sequence.
Figure 9: Ablation studies for shift magnitude $\lambda$, focused on Split CP experiments on the blue protein design dataset, where each row corresponds to a different shift magnitude $\lambda\in\{1.0,3.0,5.0,7.0\}$. All experiments are conducted with the MLPRegressor from scikit-learn (with LBFGS solver, logistic activation, and default parameters otherwise) as the ML predictor, initially trained on 32 datapoints and with 32 points in the calibration set. All values are computed over 500 random seeds. Other than $\lambda$, full hyperparameters are as in Appendix E.1, Table 2.
F.2 Ablation Study for Initial Amount of Training and Calibration Data
These ablation studies examine the effect of changing the initial number of training points ($n_{\text{train}}$) and the initial number of calibration points ($n_{\text{cal}}$), focusing on the Spit CP red fluorescent protein design case. Other than the different initial sample sizes (where $n_{\text{train}}=n_{\text{cal}}$), the experimental setting and other hyperparameters are the same as in the blue protein results reported in the main paper Figure 2, where $n_{\text{train}}=n_{\text{cal}}=32$ (full hyperparameters reported in Appendix E.1, Table 2). All values are computed over 500 random seeds.
Increasing the initial training and calibration sample size tends to hurt the coverage of baselines and increase the discrepancy in coverage between the $\{1,2,3\}$-step CP methods in these experiments. This may be because more training data improves the performance of the ML model, which can enable it to propose more extreme design sequences. This explanation is plausible because the increase in mean protein fitness values with larger initial training sets is an indication of those design procedures being more aggressive. Interval widths of the MFCS Split CP methods tend to increase with larger initial sample sizes, potentially for the same reason, to maintain coverage when the shift is more extreme.
Figure 10: Ablation studies for the initial amount of training and calibration data ($n_{\text{train}}=n_{\text{cal}}$,) focused on Split CP experiments on the red protein design dataset, where each row corresponds to a different initial sample size $n_{\text{train}}=n_{\text{cal}}\in\{16,32,48,64\}$. The ML predictor is the MLPRegressor from scikit-learn (with LBFGS solver, logistic activation, and default parameters otherwise). All values are computed over 500 random seeds. Other than $n_{\text{train}}$ and $n_{\text{cal}}$, full hyperparameters are as in Appendix E.1, Table 2.
F.3 ACI Step Size Ablation Study
Here we report sensitivity analyses on an ACI baseline’s “step size” hyperparameter that controls how strongly the method responds to recent over- or under-coverage (Gibbs & Candès, 2021) (by adjusting its empirical quantile level with the aim of having its long-run empirical coverage be close to the target coverage rate). Larger values of the ACI step size correspond to making larger adjustments to recent mistakes, and thus make the method more adaptive. These results are for the airfoil dataset, with hyperparameters as in Appendix E.2, Table 3, aside from the ACI adaptiveness and the reported number of active learning iterations (here $T=25$). In these experiments, increasing the adaptiveness (step size) of ACI tends to hurt its coverage performance, possibly due to changes to the CP calibration set causing its retroactive updates to be “out-of-date.”
Figure 11: Ablation studies for ACI baseline’s step size over values in $\{0.005,0.25,0.05\}$. All these experiments are conducted on the airfoil dataset with the GaussianProcessRegressor from scikit-learn, initially trained on 64 datapoints and with 16 points in the calibration set, and at each active learning step the queried point is added to either training or calibration with equal probability. All values are computed over 350 random seeds.
F.4 Recursion Depth Wall-Clock Runtimes
Table 4: Wall-clock runtimes for computing estimated MFCS weights (Eq. 9) with different estimation or recursion depths $d$. In these experiments, initially $n_{\text{cal}}=16$ points are in the calibration set and every queried point is added to the calibration set deterministically. Mean and standard error values are computed over 100 repeated experiments. Increasing the depth by one increases the wall-clock runtime by over an order of magnitude (for each evaluated recursion depth $d\in\{1,...,6\}$), which could quickly become intractable for larger $d$.
MFCS recursion depth $d$ Wall-clock runtime in seconds ($\pm$ std. err.)
1 1.297 (±$(\pm 0.063)\ \mathrm{e}{-6}$
2 9.595 (±$(\pm 0.042)\ \mathrm{e}{-5}$
3 1.607 (±$(\pm 0.008)\ \mathrm{e}{-3}$
4 3.194 (±$(\pm 0.036)\ \mathrm{e}{-2}$
5 6.628 (±$(\pm 0.030)\ \mathrm{e}{-1}$
6 1.444 (±$(\pm 0.001)\ \mathrm{e}{+1}$